OpenStudy (yacoub1993):

21 is 35% of what number? A. ; 40.8 B. ; 60 C. ; 62.4 D. ; 7.4

OpenStudy (mertsj):

21=.35x Solve it

OpenStudy (yacoub1993):

\[A. 0.35=21*x;40.8\] \[B. 21=0.35* x;60\] \[C. 21=0.35* x;62.4\] \[D. x=21*0.35;7.4\]

OpenStudy (mertsj):

What did you get?

OpenStudy (yacoub1993):

Those are the chooses they gave me

OpenStudy (mertsj):

Did you solve the equation I posted?

OpenStudy (mertsj):

0 21=.35x

OpenStudy (yacoub1993):

i m solving

OpenStudy (mertsj):

Good

OpenStudy (yacoub1993):

i got 57.14 as the answer

OpenStudy (yacoub1993):

sorry i redid it and it was 60

OpenStudy (yacoub1993):

i mistyped a number

OpenStudy (mertsj):